# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ 3.2.2 Attack Surface
The state space in use.
Every QKD device, be it Alice’s or Bob’s, uses a specific state space for its operation. Each state space is a component of the system’s attack surface. In Alice’s device, the state space component of the attack surface is the span of the states she transmits. For example, in the ideal BB84, Alice transmits qubits. However, implementation vulnerabilities can enlarge this space: for example, if Alice’s device suffers from the Photon Number Splitting (PNS) vulnerability, her attack surface changes from the space spanned by $\{|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}}\}$ to the space spanned by $\{|{0,0}\rangle^{\!\text{\tiny F}},|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}},|{2,0}\rangle^{\!\text{\tiny F}},|{1,1}\rangle^{\!\text{\tiny F}},|{0,2}\rangle^{\!\text{\tiny F}}\}$.
In Bob’s device, the state space component of the attack surface is the space that affects Bob’s measurement. For example, in the ideal BB84, this is again the qubit space. However, if Bob’s device suffers from the detection timing vulnerability stated in Example 2, instead of being affected by the space spanned by $\{|{1,0,0,0}\rangle^{\!\text{\tiny F}},|{0,1,0,0}\rangle^{\!\text{\tiny F}}\}$, he will be affected by the span of $\{|{1,0,0,0}\rangle^{\!\text{\tiny F}},|{0,1,0,0}\rangle^{\!\text{\tiny F}},|{0,0,1,0}\rangle^{\!\text{\tiny F}},|{0,0,0,1}\rangle^{\!\text{\tiny F}}\}$. The connection to Measurement Space Vulnerabilities (Definition 4) is clear: if a device suffers from a Measurement Space Vulnerability, the space it measures is different from the ideal space, which modifies the attack surface and usually enlarges it, as occurs in the above example.
@@
Every device interacts with its physical environment to a certain degree. Depending on how drastically a QKD device can be affected through its interaction with the environment, or how much internal information is exposed through it, that interaction can enable critical attacks on the QKD device. For example, consider the physical configuration of Alice’s or Bob’s device when they prepare to send or measure a state. If this configuration can somehow be probed by Eve, then it is part of the attack surface of the devices, and Eve can potentially use it for an attack [17]. Section 6 defines attacks based on this part of the attack surface as “Quantum Side-Channel Attacks” and includes a detailed explanation of the above example.
3.2.3 Exploits
@@
One known example is the “faked states” attack family [19]. Attacks in this family share one common property: Eve receives the state that Alice sent, measures it, and then re-sends a state that will force a specific interpretation in Bob’s system. Using our new methodology, this attack family can be decomposed to a set of vulnerabilities where fake, out-of-protocol states can force a specific interpretation, and a single, generic exploit method (which is discussed below).
Another known example is the detector efficiency mismatch attack family [20], which is a special case of the faked states attack family. Attacks in this family rely on vulnerabilities where different detectors are more/less sensitive depending on some characteristic of the input states. The generic exploit of this attack family creates attacks that manipulate this characteristic, such that whenever Eve and Bob choose different bases, Bob is less likely to detect the modified signal.
@@
Let us denote a BB84 state entering Bob’s system as $|{\psi}\rangle_{\mathrm{B}}$, and its faked counterpart as $|{\psi_{\mathrm{fake}}}\rangle_{\mathrm{B}}$. Let the vectors $|{E_{0}}\rangle_{\mathrm{E}}^{\mathrm{anc}},|{E_{1}}\rangle_{\mathrm{E}}^{\mathrm{anc}},|{E_{2}}\rangle_{\mathrm{E}}^{\mathrm{anc}},|{E_{3}}\rangle_{\mathrm{E}}^{\mathrm{anc}}$ be four orthogonal state vectors, and let $U_{\mathrm{E}}^{c}$ and $U_{\mathrm{E}}^{H}$ be Eve’s unitary operations (depending on her basis choice).
@@
$\begin{array}[]{ccc}U_{\mathrm{E}}^{c}|{0}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{0_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{0}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\\ U_{\mathrm{E}}^{c}|{1}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{1_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{1}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\end{array}$ $\begin{array}[]{ccc}U_{\mathrm{E}}^{H}|{+}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{+_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{2}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\\ U_{\mathrm{E}}^{H}|{-}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{-_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{3}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\end{array}$ (5)
@@
4 Quantum Fuzzing: Vulnerability Research for QKD
@@
4.1 Motivation: Practical Adversaries for QKD Systems
@@
The same applies when analyzing the security of QKD systems in practice: system designers and attackers might not be fully aware of physical vulnerabilities in a given implementation, making vulnerability research an important tool for identifying vulnerabilities in QKD systems.
@@
4.2 Quantum Fuzzing: Definition
The concept of fuzzing [42, 25] is a simple yet effective method to conduct vulnerability research on a device, whether classical or quantum, by probing it with various inputs (both valid and invalid) and analyzing the results: how each input affects the device, as well as the device’s output. Adapting this concept to QKD systems, we define Quantum Fuzzing in the following way:
@@ 4.2 Quantum Fuzzing: Definition
The concept of Quantum Fuzzing has previously been discussed in the context of quantum software testing [50, 51]. However, here we wish to consider it in a different context: the one of cybersecurity and vulnerability research for QKD.
4.3 Choosing Input States
@@
In classical fuzzing, the subject of fuzzing strategies has been thoroughly researched [52, 53]. Various rules and statistics are used to determine which inputs are likely to yield valuable outcomes and should thus be tested first. In addition, some classical fuzzers use the tested device’s output to improve their test cases mid-run: if a certain input triggers an interesting behavior, it may be beneficial to test other similar inputs, or otherwise to concatenate this interesting input to the next test cases.
@@
An intuitive strategy for fuzzing QKD devices is to begin with valid input states and gradually modify them: starting from simple variations in single degrees of freedom, and progressively moving to more complex modifications. The first test cases should be the valid protocol states, in order to verify that the device responds to them appropriately. Then, the next test cases should add small variations to the valid states. Good candidates for these variations can be inspired by previous, well-known attacks: shifting the frequency of the photons [54], changing their arrival time [23, 27], changing the signal intensity [55, 21], etc. If small variations do not trigger an interesting behavior in the device, the next test cases should perform more complex variations and modify more than one degree of freedom in each test case: for example, changing both the time and frequency of a photon, or testing a broader range of pulse intensities. As test cases progress, one can increase the complexity of modifications to each degree of freedom and combine modifications to more degrees of freedom. If a test case triggers an “interesting” behavior in the device, it should be documented, and potentially combined with other test cases in order to generate more interesting behaviors.
@@
4.4 Advantages and Limitations of Quantum Fuzzing
The most significant advantage of Quantum Fuzzing is its ability to reveal vulnerabilities without requiring expert knowledge: it can be applied with minimal knowledge of the target device (it only requires the attacker to know the valid input states it is supposed to handle) and can reveal the existence of vulnerabilities that would otherwise require extensive and intimate knowledge of the device to detect. For example, in Section 7 we show how the vulnerability that underlies the Bright Illumination attack [21, 56] could have been found using Quantum Fuzzing.
@@ 4.4 Advantages and Limitations of Quantum Fuzzing
These two limitations are also true for classical fuzzing. First, since it is impossible to check a program’s behavior under all possible inputs, classical fuzzing also cannot guarantee security. Second, classical fuzzing does not always reveal full details on the detected vulnerability: for example, the mere knowledge that sending a certain message caused a website to crash does not reveal which component of the website caused the crash, the specific vulnerability in the component, or how it could be further exploited.
@@
5 Reversed-Space Attacks
@@
5.1 Reversed Space Formalism: a Guiding Example
@@
In order to demonstrate the ideas of the Reversed-Space Attack, consider the photonic BB84 implementation of Section 2.2, where Alice sends (via a single pulse) a perfect qubit encoded into the polarization of a single photon. Further suppose that Bob uses a detector that cannot distinguish a single photon from a pair of photons. Then, even though a pair of photons never arrives from Alice in that single pulse (since Alice is assumed to be ideal), the quantum state describing that option could arrive at Bob from an imperfect channel or a channel controlled by Eve, hence, ought to be taken into account.
@@
We now explain the scenario through the lens of the Reversed-Space formalism. Let $\{|{j}\rangle_{\mathrm{B}}\}_{j}=\{|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}}\}$ be the computational basis for Bob’s system. Assume without loss of generality that Bob’s device is described by a unitary transformation $U_{\mathrm{B}}$, followed by a measurement in Bob’s computational basis.333 Any generalized measurement (POVM) can be described as adding an ancilla, measuring, and possibly forgetting (treating several outcomes as the same outcome), and thus is included within our formalization. The unitary transformation $U_{\mathrm{B}}$, and thus the actual measurement that Bob performs, depends on a random input of Bob, which determines whether Bob measures the computational or Hadamard basis. Namely, to measure in the computational basis, Bob sets $U_{\mathrm{B}}={I}$ and measures the result in his computational basis. To measure in the Hadamard basis, Bob lets the pulse go through optical devices such as a polarization rotator, which, on a single qubit, has the effect of performing $U_{\mathrm{B}}=H\triangleq\tfrac{1}{\sqrt{2}}\left(\begin{smallmatrix}1&\phantom{-{}}1\\ 1&-1\end{smallmatrix}\right)$; Bob then measures the resulting pulse in the computational basis. Note that $U_{\mathrm{B}}$ is well-defined by the optical device for any arriving state, not only a single-photon “qubit” pulse.
@@
As explained above, Bob actually receives a quantum space of a larger dimension, and his measurement outcomes may be richer than a single bit, as is the case when measuring a qubit. In particular, Bob can interpret his measurement outcome in multiple ways: (i) information: when the outcome indicates a specific value sent by Alice. This corresponds, for instance, to the case where Bob receives the state $|{1,0}\rangle^{\!\text{\tiny F}}$ or $|{2,0}\rangle^{\!\text{\tiny F}}$. (ii) a loss: when nothing is measured or when Alice did not send any value. This corresponds to receiving $|{0,0}\rangle^{\!\text{\tiny F}}$. (iii) invalid outcome: any other situation where the measurement outcome is considered invalid. This corresponds to the case where Bob gets the state $|{1,1}\rangle^{\!\text{\tiny F}}$ and both its detectors click.
@@
Once we identify all the relevant states $|{j}\rangle_{\mathrm{B}}$ measured by Bob, and all possible transformations taken by its device, we can apply the reversed transformation(s) $U_{\mathrm{B}}^{-1}=U_{\mathrm{B}}^{\dagger}$ on each such state $|{j}\rangle_{\mathrm{B}}$. These states, $\{U_{\mathrm{B}}^{\dagger}|{j}\rangle_{\mathrm{B}}\}$, span the space that influences Bob’s outcome. While Alice and Bob may not be aware of that enlarged space, any security analysis must assume that the attacker is fully aware of it (even if this space is not fully available to an attacker). Indeed, the fact that Bob’s equipment measures and returns a valid result for the state $|{2,0}\rangle^{\!\text{\tiny F}}$ may be unknown to Alice and Bob, but known to Eve (e.g., due to analyzing the device or due to fuzzing). When Eve designs her attack, she can consider all the parts of the space spanned by $\{U_{\mathrm{B}}^{\dagger}|{j}\rangle_{\mathrm{B}}\}$ that is available to her, and she is not limited to using only ideal qubits as Alice and Bob believe they do.
@@
We call an attack designed according to this observation a Reversed-Space Attack for a specific reason: the term “reversed” here is borrowed from the time reversal symmetry of quantum theory. The symmetry of quantum mechanics to the exchange of the prepared (pre-selected) state and the measured (post-selected) state was suggested by [57, 58], and was already used in quantum cryptography as well, see the time-reversed EPR scheme [59] for example. Interestingly, the time-reversed EPR scheme of [59] also leads to more secure protocols, named “measurement-device-independent QKD” [5].
5.2 The Reversed-Space Attack

### v3

@@ 3.2.2. Attack Surface
The state spaces in use.
Every QKD device, be it Alice’s or Bob’s, uses a specific state space for its operation. Each state space is a component of the system’s attack surface. In Alice’s device, the state space component of the attack surface is the span of the states she transmits. For example, in the ideal BB84, Alice transmits qubits. However, implementation vulnerabilities can enlarge this space: for example, if Alice’s device suffers from the Photon Number Splitting (PNS) vulnerability, her attack surface changes from the space spanned by $\{|{0{,}1}\rangle^{\!\text{\tiny F}},|{1{,}0}\rangle^{\!\text{\tiny F}}\}$ to the space spanned by $\{|{0{,}0}\rangle^{\!\text{\tiny F}},|{0{,}1}\rangle^{\!\text{\tiny F}},|{1{,}0}\rangle^{\!\text{\tiny F}},|{0{,}2}\rangle^{\!\text{\tiny F}},|{1{,}1}\rangle^{\!\text{\tiny F}},|{2{,}0}\rangle^{\!\text{\tiny F}}\}$.
In Bob’s device, the state space component of the attack surface is the space that affects Bob’s measurement. For example, in the ideal BB84, this is again the qubit space. However, if Bob’s device suffers from the detection timing vulnerability stated in Example 2, instead of being affected by the space spanned by $\{|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}},|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}\}$, he will be affected by the span of $\{|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}},|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}},|{0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}},|{1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\}$. The connection to Measurement Space Vulnerabilities (Definition 4) is clear: if a device suffers from a Measurement Space Vulnerability, the space it measures is different from the ideal space, which modifies the attack surface and usually enlarges it, as occurs in the above example.
@@
Every device interacts with its physical environment to a certain degree. Depending on how drastically a QKD device can be affected through its interaction with the environment, or how much internal information is exposed through it, that interaction can enable critical attacks on the QKD device. For example, consider the physical configuration of Alice’s or Bob’s device when they prepare to send or measure a state. If this configuration can somehow be probed by Eve, then it is part of the attack surface of the devices, and Eve can potentially use it for an attack (vakhitov_large_2001, ). Section 6 defines attacks based on this part of the attack surface as “Quantum Side-Channel Attacks” and includes a detailed explanation of the above example.
3.2.3. Exploits
@@
One known example is the “faked states” attack family (makarov_faked_2005, ). Attacks in this family share one common property: Eve receives the state that Alice sent, measures it, and then re-sends a state that will force a specific interpretation in Bob’s system. Using our new methodology, this attack family can be decomposed to a set of vulnerabilities where fake, out-of-protocol states can force a specific interpretation, and a single, generic exploit method (which is discussed below).
Another known example is the detector efficiency mismatch attack family (makarov_faked_2008, ), which is a special case of the faked states attack family. Attacks in this family rely on vulnerabilities where different detectors are more/less sensitive depending on some characteristic of the input states. The generic exploit of this attack family creates attacks that manipulate this characteristic, such that whenever Eve and Bob choose different bases, Bob is less likely to detect the modified signal.
@@
Let us denote a BB84 state entering Bob’s system as $|{\psi}\rangle_{\mathrm{B}}$, and its faked counterpart as $|{\psi_{\mathrm{fake}}}\rangle_{\mathrm{B}}$. Let the vectors $|{E_{0}}\rangle_{\mathrm{E}}^{\mathrm{anc}},|{E_{1}}\rangle_{\mathrm{E}}^{\mathrm{anc}},|{E_{2}}\rangle_{\mathrm{E}}^{\mathrm{anc}},|{E_{3}}\rangle_{\mathrm{E}}^{\mathrm{anc}}$ be four orthogonal state vectors, and let $U_{\mathrm{E}}^{\mathrm{c}}$ and $U_{\mathrm{E}}^{\mathrm{H}}$ be Eve’s unitary operations (depending on her basis choice).
@@
(5) $\begin{array}[]{ccc}U_{\mathrm{E}}^{c}|{0}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{0_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{0}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\\ U_{\mathrm{E}}^{c}|{1}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{1_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{1}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\end{array}$ $\begin{array}[]{ccc}U_{\mathrm{E}}^{H}|{+}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{+_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{2}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\\ U_{\mathrm{E}}^{H}|{-}\rangle_{\mathrm{A}}|{0}\rangle_{\mathrm{E}}^{\mathrm{anc}}&=&|{-_{\mathrm{fake}}}\rangle_{\mathrm{B}}|{E_{3}}\rangle_{\mathrm{E}}^{\mathrm{anc}}~,\end{array}$
@@
4. Quantum Fuzzing: Vulnerability Research for QKD
@@
4.1. Motivation: Practical Adversaries for QKD Systems
@@
The same applies when analyzing the security of QKD systems in practice: system designers and attackers might not be fully aware of vulnerabilities in a given implementation, making vulnerability research an important tool for identifying vulnerabilities in QKD systems.
@@
4.2. Quantum Fuzzing: Definition
The concept of fuzzing (fuzzing, ; shellcoders_handbook, ) is a simple yet effective method to conduct vulnerability research on a device, whether classical or quantum, by probing it with various inputs (both valid and invalid) and analyzing the results: how each input affects the device, as well as the device’s output. Adapting this concept to QKD systems, we define Quantum Fuzzing in the following way:
@@ 4.2. Quantum Fuzzing: Definition
The concept of Quantum Fuzzing has previously been discussed in the context of quantum software testing (quanfuzz, ; quantfuzz2, ). However, here we wish to consider it in a different context: the one of cybersecurity and vulnerability research for QKD.
4.3. Choosing Input States
@@
In classical fuzzing, the subject of fuzzing strategies has been thoroughly researched (fuzzing_review_1, ; fuzzing_review_2, ). Various rules and statistics are used to determine which inputs are likely to yield valuable outcomes and should thus be tested first. In addition, some classical fuzzers use the tested device’s output to improve their test cases mid-run: if a certain input triggers an interesting behavior, it may be beneficial to test other similar inputs, or otherwise to concatenate this interesting input to the next test cases.
@@
An intuitive strategy for fuzzing QKD devices is to begin with valid input states and gradually modify them: starting from simple variations in single degrees of freedom, and progressively moving to more complex modifications. The first test cases should be the valid protocol states, in order to verify that the device responds to them appropriately. Then, the next test cases should add small variations to the valid states. Good candidates for these variations can be inspired by previous, well-known attacks: shifting the frequency of the photons (tans_attack, ), changing their arrival time (qi_time-shift_2006, ; gelles_security_2012, ), changing the signal intensity (gottesman_security_2004, ; lydersen_blinding_commercial_2010, ), etc. If small variations do not trigger an interesting behavior in the device, the next test cases should perform more complex variations and modify more than one degree of freedom in each test case: for example, changing both the time and frequency of a photon, or testing a broader range of pulse intensities. As test cases progress, one can increase the complexity of modifications to each degree of freedom and combine modifications to more degrees of freedom. If a test case triggers an “interesting” behavior in the device, it should be documented, and potentially combined with other test cases in order to generate more interesting behaviors.
@@
4.4. Advantages and Limitations of Quantum Fuzzing
The most significant advantage of Quantum Fuzzing is its ability to reveal vulnerabilities without requiring expert knowledge: it can be applied with minimal knowledge of the target device (it only requires the attacker to know the valid input states it is supposed to handle) and can reveal the existence of vulnerabilities that would otherwise require extensive and intimate knowledge of the device to detect. For example, in Section 7 we show how the vulnerability that underlies the Bright Illumination attack (lydersen_blinding_commercial_2010, ; sague_blinding_pulsed_2011, ) could have been found using Quantum Fuzzing.
@@ 4.4. Advantages and Limitations of Quantum Fuzzing
These two limitations are also true for classical fuzzing. First, since it is impossible to test a generic program’s behavior under all possible inputs, classical fuzzing also cannot guarantee security. Second, classical fuzzing does not always reveal full details on the detected vulnerability: for example, the mere knowledge that sending a certain message caused a website to crash does not reveal which component of the website caused the crash, the specific vulnerability in the component, or how it could be further exploited.
@@
5. Reversed-Space Attacks
@@
5.1. Reversed Space Formalism: a Guiding Example
@@
Consider a photonic BB84 implementation where Alice sends (via a single pulse) a perfect qubit into the polarization of a single photon (as discussed in Section 2.2). Further suppose that Bob uses a detector that cannot distinguish a single photon from a pair of photons. Then, even though a pair of photons never arrives from Alice in that single pulse (since Alice is assumed to be ideal), the quantum state describing that option could arrive at Bob from an imperfect channel or a channel controlled by Eve, hence, ought to be taken into account.
@@
We now explain the scenario through the lens of the Reversed-Space formalism. Let $\{|{j}\rangle_{\mathrm{B}}\}_{j}=\{|{0{,}1}\rangle^{\!\text{\tiny F}},|{1{,}0}\rangle^{\!\text{\tiny F}}\}$ be the computational basis for Bob’s system. Assume without loss of generality that Bob’s device is described by a unitary transformation $U_{\mathrm{B}}$, followed by a measurement in Bob’s computational basis. [^3] Any generalized measurement (POVM) can be described as adding an ancilla, measuring, and possibly forgetting (treating several outcomes as the same outcome), and thus is included within our formalization. The unitary transformation $U_{\mathrm{B}}$, and thus the actual measurement that Bob performs, depends on a random input of Bob, which determines whether Bob measures the computational or Hadamard basis. Namely, to measure in the computational basis, Bob sets $U_{\mathrm{B}}={I}$ and measures the result in his computational basis. To measure in the Hadamard basis, Bob lets the pulse go through optical devices such as a polarization rotator, which, on a single qubit, has the effect of performing $U_{\mathrm{B}}=H\triangleq\tfrac{1}{\sqrt{2}}\left(\begin{smallmatrix}1&\phantom{-{}}1\\ 1&-1\end{smallmatrix}\right)$; Bob then measures the resulting pulse in the computational basis. Note that $U_{\mathrm{B}}$ is well-defined by the optical device for any arriving state, not only a single-photon “qubit” pulse.
@@
As explained above, Bob actually receives a quantum space of a larger dimension, and his measurement outcomes may be richer than a single bit, as is the case when measuring a qubit. In particular, Bob can interpret his measurement outcome in multiple ways: (i) information: when the outcome indicates a specific value sent by Alice. This corresponds, for instance, to the case where Bob receives the state $|{0{,}1}\rangle^{\!\text{\tiny F}}$ or $|{0{,}2}\rangle^{\!\text{\tiny F}}$. (ii) a loss: when nothing is measured or when Alice did not send any value. This corresponds to receiving $|{0{,}0}\rangle^{\!\text{\tiny F}}$. (iii) invalid outcome: any other situation where the measurement outcome is considered invalid. This corresponds to the case where Bob gets the state $|{1{,}1}\rangle^{\!\text{\tiny F}}$ and both its detectors click.
@@
Once we identify all the relevant states $|{j}\rangle_{\mathrm{B}}$ measured by Bob, and all possible transformations taken by its device, we can apply the reversed transformation(s) $U_{\mathrm{B}}^{-1}=U_{\mathrm{B}}^{\dagger}$ on each such state $|{j}\rangle_{\mathrm{B}}$. These states, $\{U_{\mathrm{B}}^{\dagger}|{j}\rangle_{\mathrm{B}}\}$, span the space that influences Bob’s outcome. While Alice and Bob may not be aware of that enlarged space, any security analysis must assume that the attacker is fully aware of it (even if this space is not fully available to an attacker). Indeed, the fact that Bob’s equipment measures and returns a valid result for the state $|{0{,}2}\rangle^{\!\text{\tiny F}}$ may be unknown to Alice and Bob, but known to Eve (e.g., due to analyzing the device or due to fuzzing). When Eve designs her attack, she can consider all the parts of the space spanned by $\{U_{\mathrm{B}}^{\dagger}|{j}\rangle_{\mathrm{B}}\}$ that is available to her, and she is not limited to using only ideal qubits as Alice and Bob believe they do.
@@
We call an attack designed according to this observation a Reversed-Space Attack for a specific reason: the term “reversed” here is borrowed from the time reversal symmetry of quantum theory. The symmetry of quantum mechanics to the exchange of the prepared (pre-selected) state and the measured (post-selected) state was suggested by (ABL64, ; AV90, ), and was already used in quantum cryptography as well, see the time-reversed EPR scheme (BHM96, ) for example. Interestingly, the time-reversed EPR scheme of (BHM96, ) also leads to more secure protocols, named “measurement-device-independent QKD” (Xu2020, ).
5.2. The Reversed-Space Attack